Number of Components Inspection Time
33 85
14 50
7 31
18 59
16 52
12 41
24 72
43 100
6 21
12 42
18 64
8 25
31 79
13 49
12 30
20 62
18 52
20 59
24 73
43 101
17 59
13 45
22 67
13 45
24 69

a-1. Estimate the linear, quadratic, and cubic regression models. Report the Adjusted R2 for each model. (Round answers to 4 decimal places.)


a-2. Which model has the best fit?


  • Linear model

  • Quadratic model

  • Cubic model


b. Use the best model to predict the time required to inspect a device with 37 components. (Round coefficient estimates to at least 4 decimal places and final answer to 2 decimal places.)
Answer #1

We use the following data to run the regression.

Inspection Time (t) Number of Components (x) X^2 X^3
85 33 1089 35937
50 14 196 2744
31 7 49 343
59 18 324 5832
52 16 256 4096
41 12 144 1728
72 24 576 13824
100 43 1849 79507
21 6 36 216
42 12 144 1728
64 18 324 5832
25 8 64 512
79 31 961 29791
49 13 169 2197
30 12 144 1728
62 20 400 8000
52 18 324 5832
59 20 400 8000
73 24 576 13824
101 43 1849 79507
59 17 289 4913
45 13 169 2197
67 22 484 10648
45 13 169 2197
69 24 576 13824

a-1) We obtain the following results -

For linear regression:

Regression Statistics
Multiple R 0.97
R Square 0.94
Adjusted R Square 0.93
Standard Error 5.40
Observations 25.00
ANOVA
df SS MS F Significance F
Regression 1.00 9849.22 9849.22 338.20 0.00
Residual 23.00 669.82 29.12
Total 24.00 10519.04
Coefficients Standard Error t Stat P-value Lower 95% Upper 95% Lower 95.0% Upper 95.0%
Intercept 17.52 2.42 7.25 0.00 12.52 22.52 12.52 22.52
Number of Components (x) 2.07 0.11 18.39 0.00 1.83 2.30 1.83 2.30

For quadratic regression -

Regression Statistics
Multiple R 0.90
R Square 0.81
Adjusted R Square 0.80
Standard Error 9.31
Observations 25.00
ANOVA
df SS MS F Significance F
Regression 1.00 8524.06 8524.06 98.27 0.00
Residual 23.00 1994.98 86.74
Total 24.00 10519.04
Coefficients Standard Error t Stat P-value Lower 95% Upper 95% Lower 95.0% Upper 95.0%
Intercept 39.56 2.58 15.33 0.00 34.22 44.90 34.22 44.90
X^2 0.04 0.00 9.91 0.00 0.03 0.05 0.03 0.05

For cubic regression -

Regression Statistics
Multiple R 0.83
R Square 0.68
Adjusted R Square 0.67
Standard Error 12.01
Observations 25.00
ANOVA
df SS MS F Significance F
Regression 1.00 7204.19 7204.19 49.99 0.00
Residual 23.00 3314.85 144.12
Total 24.00 10519.04
Coefficients Standard Error t Stat P-value Lower 95% Upper 95% Lower 95.0% Upper 95.0%
Intercept 46.59 2.84 16.42 0.00 40.72 52.46 40.72 52.46
X^3 0.00 0.00 7.07 0.00 0.00 0.00 0.00 0.00

a-2) Based on the R sqaure and adjusted r square, we see the linear model has the best fit.

a-3) To inspect a device with 37 components using the linear regression we use the following equation-

t = 17.52 + 2.07 * X = 17.52 + 2.07 * 37 = 94